# Keystroke Dynamics: Concepts, Techniques, and Applications

### v2

@@ Abstract
Reliably identifying and verifying subjects remains integral to computer system security. Various novel authentication techniques such as biometric authentication systems have been devised in recent years. This paper surveys keystroke-based authentication systems and their applications.
Keystroke dynamics is a behavioral biometric that is emerging as an important tool for cybersecurity as it promises to be non-intrusive and cost-effective. Also, no additional hardware is required, making it convenient to deploy. This survey covers novel keystroke datasets, state-of-the-art keystroke authentication algorithms, keystroke authentication on touch screen and mobile devices, and various prominent applications of such techniques beyond authentication. The paper covers all the significant aspects of keystroke dynamics and can be considered as a reference for future researchers in this domain.
The paper includes a discussion of the latest keystroke datasets, providing researchers with up-to-date resources for analysis and experimentation. Additionally, we review the state-of-the-art algorithms adopted within this domain, offering insights into the cutting-edge techniques utilized for keystroke analysis. Moreover, our paper explains the diverse applications of keystroke dynamics, particularly focusing on security, verification and identification uses. Beyond these crucial areas, we mention other additional applications where keystroke dynamics can be applied, broadening the scope of understanding regarding its potential impact across various domains. Unlike previous survey papers that typically concentrate on specific aspects of keystroke dynamics, our comprehensive analysis presents all relevant areas within this field. By introducing discussions on the latest advancements, we provide readers with a thorough understanding of the current landscape and emerging trends in keystroke dynamics research. Furthermore, our paper presents a summary of future research opportunities, highlighting potential areas for exploration and development within the realm of keystroke dynamics. This forward-looking perspective aims to inspire further inquiry and innovation, guiding the trajectory of future studies in this dynamic field.
Keywords— keystroke, authentication, behavioral, identification, imposter, detection, security.
@@ 1 INTRODUCTION
User authentication is a process which verifies that someone who is attempting to access services and applications is who they claim to be. There are three broad categories of approaches to authenticating a user. The first category is based on What you know, such as the username and password or answers to security questions. This is also known as Knowledge-based Authentication (KBA). The second category is based on What you have, i.e, a physical object that one can possess, e.g., a security badge or a security key such as Yubikey. The third category is based on What you are, which can be i) physiological, e.g., face, fingerprint, or iris ii) behavioral, e.g., keystroke dynamics, mouse dynamics, mobile motion, or swipes. While KBA currently is still the dominant approach for identity management, there are active efforts in creating alternatives such as face recognition and FIDO. In particular, an emerging field of user authentication is behavioral biometrics, which authenticates a user based on their behavior when interacting with a computing device(What you are).
Keystroke biometric authentication systems are a convenient and user friendly authentication method where a person is authenticated based on their typing pattern. Though it shows great potential, its real world application so far has been limited. Keystroke biometric authentication systems have several advantages.
@@
    They are cost effective as no additional hardware is required, only the regular keyboard is needed.
@@
    They are non-intrusive and does not create any extra hassle for the user.
@@
Keystroke authentication operates by creating a template for each user based on their typing pattern during an enrollment period. Once enrolled, the test sample from the user is contrasted with the representation of that same user and a matching score will be calculated. Matching scores are calculated based on the timing features of keystrokes. Figure 1 provides an overview of the timing features of keystrokes.
@@
Researchers use three kinds of keystroke data for authentication.
@@
    Free-text, also known as dynamic text, permits the user to type freely without any restriction [85]. For example, if a user writes a paragraph on a topic, it will be free-text. Free-text based keystroke dynamics can enable continuous authentication as a user types.
@@
    Semi fixed-text, shares some characteristics with free and fixed-text. Example of semi fixed-text include Linux commands [126] and the webform keystroke data [125].
@@ 1.1 Justification for a New Survey
The oldest survey article we reviewed is Alsultan and Warwick’s article from 2013. Alsultan and Warwick [12] focused on free-text keystroke systems. Fixed-text or semi fixed-text keystroke systems were not included in their survey. In our survey, we focus on all kinds of keystroke systems. Zhong and Deng [131] presented a comprehensive survey of keystroke dynamics; Giot, Dorizzi, and Rosenberger [43] provided a review of previously created keystroke benchmarking datasets; Ali et al. [10] presented an elaborate survey of the keystroke data collection details and classification algorithms. However, as these surveys were published in 2015, they were not able to cover more recent trends in keystroke dynamics research or updated benchmark datasets e.g., [24], [114], [85], [125], [124] etc. available since 2015.
Saini, Kaur, and Bhatia [98] and Saifan et al. [97] presented a comprehensive review on keystroke dynamics research - especially for touch screen and mobile devices. But these papers are not up to date as these were published in 2016. Also these papers did not consider all kinds of devices for keystroke dynamics research - their only focus was mobile devices.
Shinde, Shetty and Mehra [105] mainly focused on different aspects of static keystroke authentication. They did not discuss dynamic or continuous keystroke authentication in their survey. In our survey, we explain both static and dynamic keystroke authentication. Ali et al. [9] provided a comprehensive assessment of the newest datasets and algorithms used in keystroke dynamics based research. This survey was conducted in 2017 and new datasets and algorithms are now available.
Sanghi and Arya [100] presented a review of the methods and metrics used in keystroke dynamics. This paper described various applications of keystroke dynamics as well. Sadikan, Ramli, and Fudzee [95] performed a comprehensive analysis of the applications of keystroke dynamics. From this paper, we mention the latest applications of keystroke dynamics. Tewari [118] discussed only deep learning algorithms for keystroke dynamics. Conversly, in this paper we focus on many differing types of algorithms for keystroke dynamics.
In this survey, we present keystroke dynamics performance metrics, the latest benchmark datasets for keystroke dynamics, state of the art keystroke authentication algorithms, keystroke feature engineering techniques, keystroke dynamics for touch screen and mobile devices, and the most important applications of keystroke dynamics.
Table 1: A summary of recent survey papers focusing on keystroke dynamics. The surveys are compared across five categories based on whether they cover various areas viz. datasets, algorithms, mobile device keystroke capture, feature engineering approaches, and applications.
@@
Table 1 shows a comparison of existing surveys on keystroke dynamics. This paper surveys recent papers on keystroke dynamics. We highlight their unique contributions, advantages and disadvantages. We have categorized the studies based on their contribution to one of the following categories: keystroke datasets, keystroke authentication algorithms, keystroke feature engineering approaches, keystroke authentication on touch screen and mobile devices, applications of keystroke dynamics.
@@ 1.2 Methodology
Through this paper, we cover all recent developments of keystroke dynamics research and their applications. We collected papers from 2013-present on keystroke dynamics from sources such as Google Scholar, IEEE and Academia.edu. The following search terms were used:
@@
  • Any sort of novel method to improve the authentication performance of keystroke dynamics based systems such as novel/efficient keystroke authentication algorithm or software, effective keystroke feature engineering approach
@@
If a paper contributed in any of the above areas, we included/discussed that paper in our survey. Predominantly papers published after 2013 were considered for this survey as we wanted to cover the developments of keystroke dynamics during the last decade. Previous surveys on keystroke dynamics focused on specific areas such as keystroke dynamics on touch screen and mobile devices, keystroke authentication algorithms and keystroke datasets. In this paper, we cover all the areas of keystroke dynamics. An overview of the structure of this paper is shown in Figure 2.
@@ 2 Keystroke Biometrics
Keystroke dynamics research has been ongoing for decades and the main intended application has been verification and identification of subjects. Modern computer systems contain vital user information making it important to ensure the cyber security of computer systems. Authentication via something a user knows (i.e. a password) is no longer sufficient to guarantee the security of sensitive data.
In the context of keystroke dynamics, authentication, verification, and identification have different goals. Authentication means ensuring that a user is who they claim to be by comparing their typing patterns with a previously stored profile. In order to confirm the access of the correct individual, authentication is used during login to the system. Verification process can be applied to different stages of interaction. In order to detect unauthorized/imposter access continuously during a session, verification is performed. Identification process seeks to determine the identity of an unknown user from a pool of candidates based on their typing patterns. In identification process, the input is effectively matched to the most similar profile within a database.
@@ 2.1 Security and Authentication
Sridhar, Vaidya and Yawalkar [108] proposed a keystroke authentication system for intruder detection in secure virtual environments. For this purpose, authentication data from login attempts were stored. Feature sets were extracted from 200 samples of username and password data. Fuzzy Logic algorithms were used in order to develop signatures for each participant. The typing pattern of each participant was compared to the respective signature in order to authenticate the participants. Membership functions of fuzzy logic were implemented for authenticating a participant. For 200 samples, they achieved a FAR of 0.0% and a FRR 0f 0.0% for both Bell MF and Gaussian MF.
@@ 2.2 Person Identification and Forensic Investigation
Mondal and Bours [83] proposed three methods in order to identify a person from his/her typing behavior. They proposed pairwise user coupling technique along with different machine learning algorithms. They also tried to analyze person identification via an optimized feature set. For their experiment, they used a novel keystroke dataset. The dataset was gathered from three online exams. The subjects were PACE University undergraduates. 64 students participated in the data collection. The subjects typed minimum 500 keystrokes on each of the three exams. Mondal and Bours used a subset of data from the first exam as a training set. The remaining data was used for testing. They implemented Pairwise User Coupling (PUC) technique on their dataset using the machine learning classifiers. The best accuracy achieved was 89.7%.
@@
Mohlala, Ikuesan, and Venter [81] proposed a method based on keystroke dynamics in order to collect evidence for digital forensic readiness. The main goal was to collect solid keystroke evidence for forensic investigations. Their proposed model consisted of two phases - 1) the pattern development phase and 2) the pattern testing/validation phase. In the development phase, input data was collected from the subject. They used the output of this phase as the input for the next phase which was the pattern identification phase. In the data storage and preservation phase, the extracted pattern was stored in the database. A simple regression model decision process was used for subject attribution process. In the application development phase, WEKA machine learning libraries were used for feature classification. The extracted features were - key hold time, down-down time, up-down time, key press pressure, finger area, average hold time, average finger area and average pressure. They used an Android application with a software keyboard for recording raw data. The dataset consisted of 42 subjects. The subjects typed fixed password with 51 samples per subject. For the experiment process, Mohlala, Ikuesan, and Venter implemented their developed tool on the benchmark dataset [14]. In the pattern development process, 1574 rules were extracted. Each subject’s pattern was accurately attributed to the respective user on the test set, except for one user. However, all subjects were accurately attributed using the attribution mechanism. Based on the results, it was clear that this study could achieve reliable accuracy for subject attribution in digital forensics.
2.3 User Authentication for Online Assesments and Examinations
@@
In [28], Chen et al. designed a keystroke dynamics based user authentication system for online assesments and examinations. Their proposed framework was based on edge computing architecture. There were keystroke profiles of the users in the framework. In this study, Chen et al. used four features - hold time (HT), press flight time (FT), HT proportion (HTP) and FT proportion (FTP). There were two parts in the profile for static authentication and continuous authentication respectively. In order to perform static keystroke authentication, they used an anomaly detector based on Gaussian model. In the case of continuous keystroke authentication, keystroke events were collected continuously by the framework in order to form a continuous data stream of extracted keystroke timing features. For the evaluation purpose, three public datasets were used ([66], [47], [114]). In order to investigate the effectiveness of the framework, Chen et al. used a real-world case of an online examination. In order to verify their basic detector, they used this dataset [66]. Chen et al. achieved an EER of 6.62%. In order to evaluate the performance of keystroke static authentication, they used dataset [47] and achieved an EER of 5.71%. EER could be decreased to 4.03% using dynamic user profiles mechanism. In order to evaluate the performance of keystroke continuous authentication, they used dataset [114]. Chen et al. achieved an average EER of 2%. The edge computing architecture was implemented to develop a prototype of online examination system. There were three phases in the online examination system - registration phase, login phase and testing phase. In order to simulate possible security related situations, 27 users participated in the case study. The obtained results for this case study were promising.
@@ 2.4 Keystroke Dynamics Performance Metrics
There are a number of performance metrics used in keystroke dynamics and each is used to present some specific information about the biometric system. This section describes all the metrics and when they are preferably used.
@@ 2.4.1 False Accept Rate and False Reject Rate
The most common performance metric used in biometric systems are the False Accept Rate (FAR) and the False Reject Rate (FRR). The FAR is the measure of the likelihood that the biometric system will incorrectly grant access to an unauthorized user [36]. This is calculated as the ratio of the number of false acceptances divided by the total number of impostor attempts, as shown in equation 1.
@@
The ideal value for both FAR and FRR is 0%, which translates to granting access to all genuine users while all impostors are caught. However, real world biometric systems are not ideal. Therefore, a lower value for both FAR and FRR is preferred instead. A system with 0% FAR is considered to have utmost security, while a system with 0% FRR is highly convenient. There is usually a trade-off between security and convenience, as a highly secure system may reject too many genuine users, thus being inconvenient while a highly convenient system may be insecure and therefore ineffective. The FAR is also known as the False Positive Rate (FPR), while the FRR is 1 - True Positive Rate (TPR).

### v3

@@ Abstract
Reliably identifying and verifying subjects remains integral to computer system security. Various novel authentication techniques, such as biometric authentication systems, have been developed in recent years. This paper provides a detailed review of keystroke-based authentication systems and their applications. Keystroke dynamics is a behavioral biometric that is emerging as an important tool for cybersecurity as it promises to be nonintrusive and cost-effective. In addition, no additional hardware is required, making it convenient to deploy. This survey covers novel keystroke datasets, state-of-the-art keystroke authentication algorithms, keystroke authentication on touch screen and mobile devices, and various prominent applications of such techniques beyond authentication. The paper covers all the significant aspects of keystroke dynamics and can be considered a reference for future researchers in this domain.
The paper includes a discussion of the latest keystroke datasets, providing researchers with an up-to-date resource for analysis and experimentation. In addition, this survey covers the state-of-the-art algorithms adopted within this domain, offering insights into the cutting-edge techniques utilized for keystroke analysis. Moreover, this paper explains the diverse applications of keystroke dynamics, particularly focusing on security, verification, and identification uses. Beyond these crucial areas, we mention additional applications where keystroke dynamics can be applied, broadening the scope of understanding regarding its potential impact across various domains. Unlike previous survey papers, which typically concentrate on specific aspects of keystroke dynamics, our comprehensive analysis presents all relevant areas within this field. By introducing discussions on the latest advances, we provide readers with a thorough understanding of the current landscape and emerging trends in keystroke dynamics research. Furthermore, this paper presents a summary of future research opportunities, highlighting potential areas for exploration and development within the realm of keystroke dynamics. This forward-looking perspective aims to inspire further inquiry and innovation, guiding the trajectory of future studies in this dynamic field.
Keywords— keystroke dynamics, behavioral biometrics, authentication, identification, cybersecurity
@@ 1 INTRODUCTION
User authentication is a process that verifies that someone who is attempting to access services and applications is who they claim to be. There are three broad categories of approaches to authenticating a user. The first category is based on What you know, such as the username and password or answers to security questions. This is also known as Knowledge-Based Authentication (KBA). The second category is based on What you have, i.e., a physical object that can be possessed, for example, a security badge or a security key such as Yubikey. The third category is based on Who you are, which can be i) physiological, e.g., face, fingerprint, or iris, ii) behavioral, e.g., keystroke dynamics, mouse dynamics, mobile motion or swipes. Although KBA is currently still the dominant approach for identity management, there are active efforts in creating alternatives such as face recognition and FIDO [^1] FIDO stands for Fast IDentity Online. It’s a set of open authentication standards designed to make online logins more secure and user-friendly by reducing reliance on passwords.. In particular, an emerging field of user authentication is behavioral biometrics, which authenticates a user based on their behavior when interacting with a computing device.
Keystroke-based biometric authentication systems are a convenient and user-friendly authentication method in which a person is authenticated based on their typing pattern. Although such systems have great potential, their application in the real world so far has been limited. Keystroke biometric authentication systems have several advantages:
@@
    They are cost-effective as no additional hardware is required; only the regular keyboard is needed.
@@
    They are non-intrusive and do not create any extra hassle for the user.
@@
Keystroke authentication operates by creating a template for each user based on their typing pattern during an enrollment period. Once enrolled, the test sample from the user is contrasted with the representation of the same user and a matching score will be calculated. Matching scores are calculated on the basis of the timing features of keystrokes. Figure 1 provides an overview of the timing features of keystrokes.
@@
Researchers use three types of keystroke data for authentication:
@@
    Free-text, also known as dynamic text, allows the user to type freely without any restriction [85]. For example, if a user writes a paragraph on a topic, it will be free-text. Free-text based keystroke dynamics can enable continuous authentication as a user types.
@@
    Semi fixed-text shares some characteristics with free and fixed-text. Examples of semi fixed-text include Linux commands [126] and the webform keystroke data [125].
@@ 1.1 Justification for a New Survey
The oldest survey article we reviewed is Alsultan and Warwick’s article from 2013 [12] that focused on free-text keystroke systems. Fixed-text or semi fixed-text keystroke systems were not included in their survey. In our survey, we focus on all three types of keystroke systems. Zhong and Deng [131] presented a comprehensive survey of keystroke dynamics; Giot, Dorizzi, and Rosenberger [43] provided a review of previously created keystroke benchmarking datasets; Ali et al. [10] presented an elaborate survey of the keystroke data collection details and classification algorithms. However, as these surveys were published in 2015, they do not cover recent trends in keystroke dynamics research or newer benchmark datasets, e.g., [24], [114], [85], [125], [124].
Saini, Kaur, and Bhatia [98] and Saifan et al. [97] presented a comprehensive review on keystroke dynamics research, especially for touch screen and mobile devices. However, these papers do not cover recent developments as they were published in 2016. Also, these papers focused only on mobile devices.
Shinde, Shetty, and Mehra [105] mainly focused on different aspects of static-text keystroke authentication. They did not discuss dynamic or continuous keystroke authentication in their survey. In our survey, we explain both static and dynamic keystroke authentication. Ali et al. [9] provided a comprehensive assessment of the newest datasets and algorithms used in keystroke dynamics research. This survey was conducted in 2017, and new datasets and algorithms are now available.
Sanghi and Arya [100] presented a review of the methods and metrics used in keystroke dynamics. This paper described various applications of keystroke dynamics as well. Sadikan, Ramli, and Fudzee [95] performed a comprehensive analysis of the applications of keystroke dynamics. In this paper, we describe the latest applications of keystroke dynamics. Tewari [118] discussed only deep learning algorithms for keystroke dynamics. This paper provides a more comprehensive description of various types of keystroke dynamics algorithms.
To summarize, this survey covers keystroke dynamics performance metrics, the latest benchmark datasets for keystroke dynamics, state-of-the-art keystroke authentication algorithms, keystroke feature engineering techniques, keystroke dynamics for touch screen and mobile devices, and the most important applications of keystroke dynamics.
Table 1: A summary of recent survey papers focusing on keystroke dynamics. The surveys are compared across five categories based on whether they cover various areas, viz. datasets, algorithms, mobile device keystroke capture, feature engineering approaches, and applications.
@@
Table 1 shows a comparison of existing surveys on keystroke dynamics. We highlight their unique contributions, advantages, and disadvantages. We have categorized the studies based on their contribution to one of the following categories: keystroke datasets, keystroke authentication algorithms, keystroke feature engineering approaches, keystroke authentication on touch screen and mobile devices, and applications of keystroke dynamics.
@@ 1.2 Methodology
Through this paper, we comprehensively cover recent developments in keystroke dynamics research and their applications. We collected papers from 2013-present on keystroke dynamics from sources such as Google Scholar, IEEE, and Academia.edu. The following search terms were used:
@@
  • New methods to improve the authentication performance of keystroke dynamics-based systems such as novel/efficient keystroke authentication algorithm or software, effective keystroke feature engineering approach
@@
If a paper contributed in any of the above areas, we have included it in this survey. Predominantly, papers published after 2013 were considered for this survey as we wanted to cover the developments of keystroke dynamics during the last decade. Previous surveys on keystroke dynamics focused on specific areas such as keystroke dynamics on touch screen and mobile devices, keystroke authentication algorithms, and keystroke datasets. An overview of the structure of this paper is shown in Figure 2.
@@ 2 Keystroke Biometrics
Keystroke dynamics research has been ongoing for decades and the main intended application has been verification and identification of subjects. Modern computer systems contain vital user information making it important to ensure the security of computer systems. Authentication via something a user knows (i.e. a password) is no longer sufficient to guarantee the security of sensitive data.
In the context of keystroke dynamics, authentication, verification, and identification have different goals. Authentication means ensuring that a user is who they claim to be by comparing their typing patterns with a previously stored profile. In order to confirm the access of the correct individual, authentication is used when logging into a system. Verification can be done during different stages of interaction. In order to detect unauthorized/imposter access continuously during a session, verification is performed. Identification seeks to determine the identity of an unknown user from a pool of candidates based on their typing patterns. In the identification process, the user input is effectively matched to the most similar profile within a database.
@@ 2.1 Security and Authentication
Sridhar, Vaidya and Yawalkar [108] proposed a keystroke authentication system for intruder detection in secure virtual environments. For this purpose, authentication data from login attempts was stored. Feature sets were extracted from 200 samples of username and password data. Fuzzy Logic algorithms were used in order to develop signatures for each participant. The typing pattern of each participant was compared to the respective signature in order to authenticate the participants. Membership functions of fuzzy logic were implemented for authenticating a participant. For 200 samples, they achieved a FAR of 0.0% and a FRR of 0.0% for both Bell MF and Gaussian MF.
@@ 2.2 Person Identification and Forensic Investigation
Mondal and Bours [83] investigated the usefulness of keystroke dynamics in identifying a person. They proposed a pairwise user coupling (PUC) technique along with different machine learning algorithms. They also analyzed person identification via an optimized feature set. For their experiment, they used a novel keystroke dataset collected from three online exams. The subjects were PACE University undergraduates. 64 students participated in the data collection. The subjects typed a minimum of 500 keystrokes in each of the three exams. Mondal and Bours used a subset of data from the first exam as a training set. The remaining data was used for testing. They implemented the PUC technique on their dataset using machine learning classifiers. The best accuracy achieved was 89.7%.
@@
Mohlala, Ikuesan, and Venter [81] proposed a method based on keystroke dynamics to collect evidence for digital forensic readiness. The main goal was to collect solid keystroke evidence for forensic investigations. Their model consisted of two core phases: 1) the pattern development phase and 2) the pattern testing/validation phase. In the development phase, input data is taken from the known user and continually developed by applying the data preprocessing, machine learning, and pattern extraction subprocesses before storing the data in the database. They used an Android application with a software keyboard to record raw data. The dataset consisted of 42 subjects. The subjects typed fixed password with 51 samples per subject. For the experiment process, Mohlala, Ikuesan, and Venter implemented their developed tool on the benchmark dataset [14]. Each subject’s pattern was accurately attributed to the respective user on the test set, except for one user. However, all subjects were accurately attributed using the attribution mechanism. Based on the results, it was clear that this study could achieve reliable accuracy for subject attribution in digital forensics.
2.3 User Authentication for Online Assessments and Examinations
@@
Chen et al. [28] designed a keystroke dynamics-based user authentication system for online assessments and examinations. Their framework is based on an edge computing architecture. There were keystroke profiles of the users in the framework. In this study, Chen et al. used four features: hold time (HT), press flight time (FT), HT proportion (HTP), and FT proportion (FTP). There were two parts in the profile for static authentication and continuous authentication respectively. In order to perform static keystroke authentication, they used an anomaly detector based on the Gaussian model. In the case of continuous keystroke authentication, keystroke events were collected continuously by the framework in order to form a continuous data stream of extracted keystroke timing features. For evaluation, three public datasets were used ([66], [47], [114]). To investigate the effectiveness of the framework, they used a real-world case of an online examination. In order to verify their basic detector, they used the dataset from Killourhy and Maxion [66] and achieved an EER of 6.62%. To evaluate the performance of static keystroke authentication, they used the dataset from Giot et al. [47] and achieved an EER of 5.71%. EER could be decreased to 4.03% using a dynamic user profiles mechanism. To evaluate the performance of continuous keystroke authentication, they used the dataset from Sun et al. [114] and achieved an average EER of 2%. The edge computing architecture was implemented to develop a prototype of an online examination system. There were three phases in the online examination system: registration phase, login phase, and testing phase. In order to simulate possible security-related situations, 27 users participated in the case study. The obtained results for this case study were promising.
@@ 2.4 Keystroke Dynamics Performance Metrics
There are a number of performance metrics used in keystroke dynamics, and each is used to present some specific information about the biometric system. This section describes all the metrics and when they are preferably used.
@@ 2.4.1 False Accept Rate and False Reject Rate
The most common performance metrics used in biometric systems are the False Accept Rate (FAR) and the False Reject Rate (FRR). The FAR is the measure of the likelihood that the biometric system will incorrectly grant access to an unauthorized user [36]. This is calculated as the ratio of the number of false acceptances divided by the total number of impostor attempts, as shown in equation 1.
@@
The ideal value for both FAR and FRR is 0%, which translates to granting access to all genuine users while all impostors are caught. However, real-world biometric systems are not ideal. Therefore, for both FAR and FRR, a value as close to zero as possible is desired. A system with 0% FAR is considered to have the utmost security, while a system with 0% FRR is highly convenient. There is usually a trade-off between security and convenience, as a highly secure system may reject too many genuine users, thus being inconvenient, while a highly convenient system may be insecure and therefore ineffective. The FAR is also known as the False Positive Rate (FPR), while the FRR = 1 - True Positive Rate (TPR).